What is the measures of two complementary angles such that the larger angle is six less than twice the smaller angle?
My solution is:
Complementary angle= 90
2x-6=90
2x=90+6
2x=96
1/2(2x)=(96)1/2
x=48
Small Complementary Angle = 48
Larger Complementary Angle = 96

Answer by lwsshak3(11628)

What is the measures of two complementary angles such that the larger angle is six less than twice the smaller angle?
**
let x=larger of two complementary angles
90-x=smaller of two complementary angles
..
x=2(90-x)-6
x=180-2x-6
3x=174
x=58
90-x=32
larger of two complementary angles=58º
smaller of two complementary angles=32º